# FaaF: Facts as a Function for the evaluation of generated text

### v2

@@ 5 Results
In the case of gpt-4-turbo and gpt-3.5-turbo, the high error rates are attributed to an overall overestimation of the factual truthfulness of a fact statement given a reference text which does not support it. Regarding claude-3-opus and claude-3-sonnet, the exceedingly high ER is primarily due to the erroneous parsing of the LM’s response. In detail, on many occasions the verification response from Claude LMs contains both True and False words, which exposes the fragility of word-matching as a means of parsing the LM’s response. Although qualitative evaluation found Claude models to be less prone to overestimation than GPT LMs, the increased verbosity of the responses some times includes phrases like ”To determine if the claim is true or false based on the given passage…”, ”…cannot be determined as true or false.” or ”…given the lack of information in the passage, we cannot determine whether the claim is True or False.” which mislead the interpretation of the response when using word matching.
@@
Asking for citations helps in the presence of correct and clear information but can also lead to false positives otherwise. The positive impact of adding citations is most evident in the ground truth answer category where the provided text always contain the required evidence to support the facts. In this instance, asking for text evidence from LMeval results in avoiding some false negative verifications. The beneficial mechanism is associated with inserting a preliminary step to the fact verification process—of the explicit use of evidence from the input text. This aligns with the findings reported in Wei et al. (2023) regarding the chain-of-thought method.
@@
Claude LMs are more reliable than GPT in correctly formatting the response for function calling. The capacity of the LM to return a correctly formatted response for function calling is distinct to their ability for accurate fact verification. In this work we allowed only one chance to respond correctly (regarding format) for all experiments and we report the cases of an failed responses for all LMs in the N/A columns in Table 2. The ER and F1micro metrics are calculated considering only the correctly formatted LM responses.
@@
As evident in Table 2, the XML format used in Claude LMs results to more reliable response formatting compared to JSON format used by GPT models. Claude LMs returned a correctly formatted response 100% of the time when using FaaF whereas GPT LMs produced some failed attempts. The failure mechanism in GPT models appears to be directly related to the citations mode—all formatting failures are seen in FaaF+citation formulations (Table 2). Looking more closely, we find that when the citation of a fact is null, there is a risk that the LM will return null in the fact verification argument as well (but only True or False are accepted according to the type annotations of the function object definition), which results to failed invocation of the FaaF function object. In the case of gpt-4-turbo, the above failure mechanism is much less pronounced compared to gpt-3.5-turbo, which shows up to 25% failure rate (71/281) in returning a correctly formatted response in the poor answers, where the citations are expected to be null most of the time.
We attempted to include mistral-large LM in our study but the failure rates associated the response formatting was prohibitive (80% in some cases). It is worth noting that the failure mechanisms of mistral-large were different to what we observed in GPT LMs. In detail, on many occasions mistral-large failed to return all the expected arguments in its response and in other cases, in stead of returning the expected JSON with the function arguments, it responded with a long-form text, which included the function arguments as part of the natural language response (which led to failure when parsing).
@@
FaaF requires less than one-fifth of the calls to the LM to perform fact verification compared to prompting. This corresponds to the average number of facts which are examined for each text (answer type) in WikiEvalFacts. Thus, the degree of efficiency improvement using FaaF is proportional to the number of facts we can encapsulate in the function object, there by avoiding their individual verification.
Table 3 presents the token counts for the tested verification formulations and model families. The token counts include tokens used for all the answer types in WikiEvalFacts dataset in each scenario. Only one model from the GPT and Claude families is included since the token count differences between models in the same family are not significant. Lastly, it should be noted that although we can constrain the completion tokens from the ML parameters we chose not to do so, to examine the token usage in the unconstrained scenario.
GPT LMs using JSON format are significantly more efficient than Claude with XML in token usage. Considering the differences between gpt-4-turbo and claude-3-opus, the observed increase in prompt and completion tokens in the FaaF approaches is associated to the tags used in the XML format which is expected from claude function calling (versus the more succinct JSON format that is expected by gpt). In the case of prompt tokens, the one-shot system prompt (which also contains XML tags) used in claude-3-opus further contributes to the token count. A respective improvement in speed is also noted, with GPT LMs being faster than Claude.
Focusing at the completion tokens, the sharp increase (more than 4X) between gpt-4-turbo and claude-3-opus seen in the case of prompt-based verification is attributed to the extra verbosity by claude-3-opus. When prompted to verify a single fact, gpt-4-turbo most times returns small phrase containing True/False. Conversely, claude-3-opus always returns a coherent explanation to justify the verification. It should be noted that the same prompt was used in both models.
@@
FaaF significantly reduces token usage, requiring fewer than one-fifth the completion tokens needed for prompt-based verification. Following the trend seen in the required LM calls, the most notable reduction in token count is achieved when replacing prompt-based verification with FaaF(T/F). In this scenario, with gpt-4-turbo half of the prompt tokens and less than one-fifth of the completion tokens are required and with claude-3-opus, two thirds of the prompt tokens and less than one-sixth of of the completion tokens are required for fact verification for a complete evaluation on WikiEval(all answer types). Including citations and the response option Not clear from the given passage progressively increases the token count due to the additional information we include in the LM verification. However, in most cases, it still remains below the token requirements of prompt-based fact verification.
@@ 6 Conclusions & future work
We show that prompt-based fact verification is prone to overestimating the truthfulness of fact statements in texts with inaccurate and/or missing information. In particular the error rate of prompting can exceed 70% when the text under review is significantly lacking in information. In such challenging situations, presenting the facts as a function (FaaF) to the language model significantly enhances the its ability to verify facts accurately. The improvement comes from leveraging a more structured generation mode of the language model. This is achieved by generating type-annotated function call arguments instead of natural language text. Additionally, it avoids the unreliable method of exact word matching to parse verification responses, a common issue in prompt-based verification methods.
@@
We examine the latest commercial language models and find significant improvement in fact verification accuracy between the larger LMs (gpt-4-turbo and claude-3-opus) and their smaller and faster counterparts (gpt-3.5-turbo and claude-3-sonnet).
GPT models using JSON function representation show a sensitivity in returning a correctly formatted response for function calling when asking from the LM to return citations to support fact verification. In contrast, Claude models using XML function representation are more reliable in this regard, consistently delivering correctly formatted responses.
Language model claude-3-opus using XML slightly outperformed gpt-4-turbo using JSON in fact verification accuracy using FaaF but required significantly more tokens to achieve that due to the more verbose format of XML tags compared to JSON.
@@
While the advantages of using function calls for fact verification are significant, there is certain variance expected in the results due to the stochastic nature of LM generation. Further extensive testing is necessary to solidify the results presented, especially probing into the influence of using XML or JSON as the means of passing the function to the LM in the performance of verification using function objects. Although, the WikiEval dataset has highlighted the importance of testing fact verification in challenging conditions and provided a convenient way to compare fact verification performance across various text qualities, it is relatively small, comprising only 50 question/answer pairs.
Additionally, the results shown with FaaF are sensitive to the instructions passed in the function object’s metadata, which highlights the need for additional research and optimisation of FaaF configurations and the interplay of the function argument’s metadata. One example of such interplay is the observation that in GPT models using JSON, including arguments for citation in the FaaF object, caused a rise in the type validation errors due to the LM confusing the different type annotations of different arguments. Other open questions include the maximum number of fact statements and the maximum permissible length for a fact that can be incorporated into a function object and whether token count is the sole limitation or if there are performance implications as well.

### v3

@@ 5 Results
In the case of GPT LMs, the high error rates are attributed to an overall overestimation of the factual truthfulness of a fact statement given a reference text which does not support it. Regarding Claude LMs, the exceedingly high ER is primarily due to the erroneous parsing of the LM’s response. In detail, on many occasions the verification response from Claude LMs contains both True and False words, which exposes the fragility of word-matching as a means of parsing the LM’s response. For example, Claude responses frequently include phrases like ”To determine if the claim is true or false based on the given passage…” which mislead the interpretation of the response when using word matching.
@@
Asking for citations helps in the presence of correct and clear information but can also lead to false positives otherwise. The positive impact of adding citations is most evident in the ground truth answer category where the provided text always contain the required evidence to support the facts. In this instance, asking for text evidence from LMeval results in avoiding some false negative verifications. The beneficial mechanism is associated with inserting a complementary step to the verification process—of the explicit use of evidence from the input text. This aligns with the findings reported in the work of Wei et al. (2023) regarding the chain-of-thought method.
@@
Claude LMs are more reliable than GPT in correctly formatting the response for function calling. As evident in the N/A values in Table 2, the XML format used in Claude LMs results to more reliable response formatting compared to JSON format used by GPT models. Claude LMs returned a correctly formatted response 100% of the time when using FaaF whereas GPT LMs produced some failed attempts. The failure mechanism in GPT models appears to be directly related to the citations mode—all formatting failures are seen in FaaF+citation formulations (Table 2). Looking more closely, we find that when the citation of a fact is null, there is a risk that the LM will return null in the fact verification argument as well (but only True or False are accepted according to the type annotations of the function object definition), which results to failed invocation of the FaaF function object.
@@
FaaF requires less than one-fifth of the calls to the LM compared to prompting. As seen in Figure 3, this corresponds to the average number of facts which are examined for each text (answer type) in WikiEvalFacts. Thus, the degree of efficiency improvement using FaaF is proportional to the number of facts we can encapsulate in the function object.
@@
A similar five-fold reduction can be seen in completion tokes usage by replacing prompting with FaaF (Table 3). Including citations and the response option Not clear from the given passage progressively increases the token count due to the additional information we include in the LM verification. However, in most cases, it still remains below the token requirements of prompt-based fact verification.
GPT using JSON format are significantly more efficient than Claude with XML in token usage. Considering the differences between gpt-4-turbo and claude-3-opus, the observed increase in prompt and completion tokens in the FaaF approaches is associated to the tags used in the XML format which is expected from claude function calling (versus the more succinct JSON format that is expected by GPT). A respective improvement in speed is also noted, with GPT LMs being faster than Claude.
Focusing at the completion tokens, the sharp increase (more than 4X) between gpt-4-turbo and claude-3-opus seen in the case of prompt-based verification is attributed to the extra verbosity by claude-3-opus.
@@ 6 Conclusions & future work
We show that prompt-based fact verification is prone to overestimating the truthfulness of fact statements in texts with inaccurate and/or missing information. In such challenging situations, presenting the facts as a function (FaaF) to the language model significantly enhances the its ability to verify facts accurately. The improvement comes from leveraging a more structured generation mode of the language model and avoiding exact word matching to interpret the LM’s response.
@@
GPT models using JSON function representation are more sensitive to formatting errors than Claude models using XML but XML is significantly more expensive token-wise.
@@
While the advantages of using function calls for fact verification are significant, further extensive testing is necessary to solidify the results presented. Although, the WikiEval dataset has highlighted the importance of testing fact verification in challenging conditions and provided a convenient way to compare fact verification performance across various text qualities, it is relatively small, comprising only 50 question/answer pairs.
Additionally, the results shown with FaaF are sensitive to the instructions passed in the function object’s metadata, which highlights the need for additional research and optimisation of FaaF configurations and the interplay of the function argument’s metadata.
Other open questions include the maximum number of fact statements and the maximum permissible length for a fact that can be incorporated into a function object and whether token count is the sole limitation or if there are performance implications as well.